OpenStudy (anonymous):

Find the slope of the line 4x+5y=28

OpenStudy (anonymous):

m=-0.8

OpenStudy (anonymous):

Write the equation of the line that satisfies the given condition: contains the point (2,6) and is perpendicular to the line 4x-5y=7. First, put the equation in slope intercept form:highlight%28y=mx%2Bb%29, where m=slope and b=y-intercept. 4x-5y=7 -4x%2B4x-5y=-4x%2B7 -5y=-4x%2B7 -5y%2F-5y=%28-4%2F-5%29x-7%2F5 y=%284%2F5%29x-7%2F5 y=highlight%284%2F5%29x-7%2F5 The slope of their line is 4/5.